# Learning (Approximately) Equivariant Networks via Constrained Optimization

### v1

@@ 1 Introduction
Equivariant neural networks (NNs) [1, 2, 3, 4] leverage known symmetries to improve sample efficiency and generalization, and are widely used in computer vision [1, 3, 5, 6, 7], graph learning [8, 9, 10, 11], and 3D structure modeling [12, 13, 14, 15, 16, 17]. Despite their advantages, training these networks can be challenging. Indeed, their inherent symmetries give rise to complex loss landscapes [18] even when the model matches the underlying data symmetry [18, 19]. Additionally, real-world datasets frequently include noise, measurement bias, and other symmetry-breaking effects such as dynamical phase transitions or polar fluids [20, 21, 22, 23, 24] that undermine the assumptions of strictly equivariant models [25, 26, 27]. These factors complicate the optimization of equivariant NNs, requiring careful hyperparameter tuning and often resulting in suboptimal performance.
@@
A well-established strategy for approaching difficult optimization problems is to use homotopy (or continuation), i.e., to first solve a simple, surrogate problem and then track its solutions through a sequence of progressively harder problems until reaching the original one [28, 29]. Continuation has been used for compressive sensing [30, 31, 32] and other non-convex optimization problems [33, 34, 35], where it is closely related to simulated annealing strategies [36, 37, 38, 39, 40]. These approaches suggest that introducing equivariance gradually can facilitate training by guiding the model through a series of less stringent intermediate optimization problems. However, implementations of this strategy based on modifying the training objective [18] or the underlying model [41] require domain-specific knowledge to design the intermediate stages and craft loss penalties, limiting practical usability.
@@
This work, therefore, addresses the following question: is there an automated, general-purpose strategy to navigate these intermediate problems that precludes manual tuning and that works effectively for various models and tasks? To this end, we leverage the connection between homotopy approaches and dual methods in optimization, which both provide rigorous frameworks for solving constrained problems by addressing a sequence of relaxed problems with progressively stricter penalties. By casting the training process as a constrained optimization problem, we introduce Adaptive Constrained Equivariance (ACE), a principled framework for automatically relaxing/tightening equivariance in NNs. Our method balances downstream performance (e.g., accuracy) with equivariance, relaxing equivariance at initial stages while gradually transitioning to a strictly equivariant model, as illustrated in Fig. 1. This transition occurs in an automated, data-driven fashion, eliminating much of the trial-and-error associated with manual tuning.
@@ 2 Related Work
Equivariant NNs encode group symmetries directly in their architecture. In computer vision, for instance, equivariant convolutional NNs accomplish this by tying filter weights for different rotations and reflections [1, 2, 3, 4]. Equivariant NNs have also been extended to 3D graphs and molecular systems, as is the case of SchNet, that operates directly on atomic coordinates while remaining $\mathsf{E(3)}$-invariant [12], and Tensor Field Networks or $\mathsf{SE(3)}$ Transformers, that extend this to full roto-translation equivariance using spherical harmonics and attention, respectively [13, 42].
Scalable, parameter-efficient architectures that remain equivariant include EGNNs [43], SEGNNs [44], and methods such as Vector Neurons [45], which lift scalars to 3D vectors, achieving exact $\mathsf{SO}(3)$ equivariance in common layers with minimal extra parameters. Moreover, recent works model 3D trajectories rather than static snapshots, such as EGNO, which uses $\mathsf{SE(3)}$ equivariant layers combined with Fourier temporal convolutions to predict spatial trajectories [46].
While exact equivariance can improve sample efficiency [47, 48, 49], it may render the optimization harder, particularly when the data violates symmetry (even if partially) due to noise or other symmetry-breaking phenomena [22, 23, 24]. This has motivated the use of models with approximate or relaxed equivariance based on architectures that soften weight-sharing, such as Residual Pathway Priors [50] that adds a non-equivariant branch alongside the equivariant one, or works such as  Wang et al. [26], where weighted filter banks are used to interpolate between equivariant kernels. Recent works have formalized the trade-offs that appear due to these symmetry violations, establishing generalization bounds that specify when and how much equivariance should be relaxed [19].
Beyond architectural modifications, another line of work relaxes symmetry during training. REMUL [18] adds a penalty for equivariance violations to the training objective, adapting the penalty weights throughout training to balance accuracy and equivariance. This procedure does not guarantee that the final solution is equivariant or even the degree of symmetry achieved. An alternative approach modifies the underlying architecture by perturbing each equivariant layer and gradually reducing the perturbation during training according to a user-defined schedule [41]. This guided relaxation facilitates training and guarantees that the final solution is equivariant, but is sensitive to the schedule. Additional equivariance penalties based on Lie derivative are used to mitigate this issue, increasing the number of hyperparameters and restricting the generality of the method.
@@ 3 Training Equivariant Models
Equivariant NNs improve sample efficiency by encoding transformations from a symmetry group $G$ [47, 48, 49]. Strictly enforcing equivariance, however, limits the expressiveness of the model and create intricate loss landscapes, which can slow down optimization and hinder their training [19, 18], particularly in tasks involving partial or imperfect symmetries [20, 21, 22, 23, 24].
To temper these trade-offs, recent work relaxes symmetry during training by modifying either the optimization objective or the model itself [41, 18, 50, 26, 19]. Explicitly, let $f_{\theta,\gamma}:X\to Y$ be a NN now parameterized by both $\theta$, the network weights, and $\gamma\in\mathbb{R}^{L}$, a set of homotopic parameters representing the model non-equivariance. The model is therefore equivariant for $\gamma=0$, i.e., $f_{\theta,0}$ satisfies (1), but can behave arbitrarily if $|\gamma_{i}|>0$ for any $i$. Consider training this NN by solving
@@
where $(x_{n},y_{n})$ are samples from a data distribution $\mathfrak{D}$, $\ell_{0}$ denotes the top-line objective function (typically, mean-squared error or cross-entropy), $\ell_{\text{eq}}$ denotes an equivariance metric, and $\alpha,\beta>0$ are hyperparameters controlling their contributions to the training loss. Hence, traditional equivariant NN training amounts to using $\alpha=1$ and $\beta=\gamma=0$ in (2). In [18], a non-equivariant model ($\gamma=1$) is trained using $\ell_{\text{eq}}(f)=\sum_{n}\mathbb{E}_{g_{n}}\bigl{[}\|f\big{(}\rho_{X}(g_{n})x_{n}\big{)}-\rho_{Y}(g_{n})f(x_{n})\|^{2}\bigr{]}$ with $g_{n}$ sampled uniformly from $G$. The values of $\alpha,\beta$ are adapted to balance the contributions of each loss, i.e., to balance accuracy and equivariance. In contrast, [41] uses an equivariance metric $\ell_{\text{eq}}$ based on Lie derivatives [51, 52], fix $\alpha,\beta$, and manually decrease $\gamma$ throughout training until it vanishes, i.e., until the model is equivariant.
While effective in particular instances, these approaches are based on sensitive hyperparameters, some of which are application-dependent. This is particularly critical in [18], where both performance and equivariance depend on the values of $\alpha,\beta$. While [41] guarantees equivariance by driving $\gamma$ to zero, its performance is substantially impacted by the rate at which it vanishes. It therefore continues to use $\ell_{\text{eq}}$ even though it is not needed to ensure equivariance. What is more, it is once again susceptible to symmetry-breaking effects in the data since it strictly imposes equivariance. In the sequel, we address these challenges by doing away with hand-designed penalties ($\beta=0$) and manually-tuned schedules for $\gamma$. Before proceeding, however, we briefly comment on the implementation of $f_{\theta,\gamma}$.
@@
Throughout this paper, we consider an architecture similar to [41] built by composing equivariant layers linearly perturbed by non-equivariant components, e.g., a linear layer or a small multilayer perceptron (MLP). Formally, let
@@ 4.1 Strictly equivariant data: Equality constraints
where $(x_{n},y_{n})$ are i.i.d. samples from $\mathfrak{D}$, $\lambda\in\mathbb{R}^{L}$ collects the dual variables [53, 54] $\lambda_{i}$, and $\hat{L}$ is called the empirical Lagrangian [55, 56]. In general, solutions of (PI) and (DI) are not related due to (i) non-convexity, which hinders strong duality [57], and (ii) the empirical approximation of the expectation in (PI) by samples. Yet, for rich enough parametrizations $f_{\theta,\gamma}$, constrained learning theory shows that solutions of (DI) do generalize to those of (PI[53, 54, 55]. In other words, (DI) approximates the solution of the equivariant NN training problem (PI) without constraining the model to be equivariant.
More precisely, (DI) gives rise to a gradient descent (on $\theta,\gamma$) and ascent (on $\lambda$) algorithm (Alg. 1) that acts as an annealing mechanism adapted to the downstream task, taking into account the effect of enforcing equivariance on the objective function. Indeed, by taking $\gamma_{i}^{(0)}=1$, Alg. 1 starts by training a flexible, non-equivariant model (step 3). The gradient step on the homotopy parameters $\gamma_{i}$ (step 4) then simultaneously seeks to reduce the objective function $\ell_{0}$ and bring $\gamma_{i}$ closer to zero, targeting the constraint in (PI). The strength of that bias toward zero (namely, $\lambda_{i}$) depends on the accumulated constraint violations (step 5): the longer $\gamma_{i}$ is strictly positive (negative), the larger the magnitude of $\lambda_{i}$. Hence, the $\gamma_{i}$ can increase to expand the flexibility of the model or decrease to exploit the symmetries in the data. The initial non-equivariant NN is therefore turned into an equivariant one by explicitly taking into account the effects this has on the downstream performance, instead of relying on hand-tuned, ad hoc penalties or schedules as in [18, 41].
Naturally, the iterates $\gamma_{i}^{(t)}$ do not vanish in any finite number of iterations. In fact, the behavior of saddle-point algorithms such as Alg. 1 are intricate even in convex settings [58, 59] and its iterates may, even for small learning rates $\eta_{p},\eta_{d}$, (i) asymptotically approach zero or (ii) display oscillations that do not subside. In case (i), if we stop Alg. 1 at iteration $T$ and deploy its solution with $\gamma=0$ (i.e., $f_{\theta^{(T)},0}$), we introduce an error that depends on the magnitude of $\gamma_{i}^{(T)}$. The following theorem bounds this approximation error for the architecture in (3) under mild assumptions on its components. In particular, note that Ass. 1 holds, e.g., for MLPs with ReLU activations and weight matrices of bounded spectral norms [60].
@@
A proof of this as well as a more refined bound is available in the Section A.1. Theorem 4.1 shows that as long as the $\{\gamma_{i}^{(T)}\}$ are small enough, the error incurred from setting them to zero is negligible. This is not the case, however, when the $\gamma_{i}$ oscillate without subsiding [case (ii)]. As we argue next, this effect can be used to identify and correct for symmetry deviations in the data.
@@
2:$J_{0}^{(t)}=\dfrac{1}{N}\sum_{n=1}^{N}\ell_{0}\big{(}f_{\theta^{(t)},\gamma^{(t)}}(x_{n}),y_{n}\big{)}$
3:$\theta^{(t+1)}=\theta^{(t)}-\eta_{p}\nabla_{\theta}J_{0}^{(t)}$
4:$\gamma_{i}^{(t+1)}=\gamma_{i}^{(t)}-\eta_{p}\Bigl{(}\nabla_{\gamma_{i}}J_{0}^{(t)}+\lambda_{i}^{(t)}\Bigr{)}$ $u_{i}^{(t+1)}=u_{i}^{(t)}+\eta_{p}(\rho u_{i}^{(t)}-\lambda_{i}^{(t)})$
5:$\lambda_{i}^{(t+1)}=\lambda_{i}^{(t)}+\eta_{d}\gamma_{i}^{(t)}\vphantom{\Big{]}_{+}}$
@@
2:$J_{0}^{(t)}=\dfrac{1}{N}\sum_{n=1}^{N}\ell_{0}\big{(}f_{\theta^{(t)},\gamma^{(t)}}(x_{n}),y_{n}\big{)}$
3:$\theta^{(t+1)}=\theta^{(t)}-\eta_{p}\nabla_{\theta}J_{0}^{(t)}$
4:$\gamma_{i}^{(t+1)}=\gamma_{i}^{(t)}-\eta_{p}\Bigl{(}\nabla_{\gamma_{i}}J_{0}^{(t)}+\lambda_{i}^{(t)}\Bigr{)}$
5:$u_{i}^{(t+1)}=u_{i}^{(t)}+\eta_{p}(\rho u_{i}^{(t)}-\lambda_{i}^{(t)})$
6:$\lambda_{i}^{(t+1)}=\Big{[}\lambda_{i}^{(t)}+\eta_{d}(|\gamma_{i}^{(t)}|-u_{i}^{(t)})\Big{]}_{+}$
@@ 4.2 Partially equivariant data: Resilient constraints
In some practical scenarios, the data used for training may only partially satisfy the equivariance relation (1) due to the presence of noise, measurement bias, or other symmetry-breaking effects [20, 21, 22, 23, 24, 25, 26, 27]. In such cases, the equivariance constraints in (PI) become too stringent and the gradient of the objective with respect to $\gamma_{i}$ in step 4 of Alg. 1 does not vanish as $\gamma_{i}$ approaches zero. Alg. 1 will then oscillate as steps 4 and 5 push $\gamma_{i}$ alternatingly closer and further from zero [case (ii)] or $\gamma_{i}$ will settle, but $\lambda_{i}$ will not vanish. These observations can then be used as evidences that the data does not fully adhere to imposed symmetries. This is not necessarily a sign that equivariance should be altogether abandoned, but that we can improve performance by relaxing it. To this end, we rely on the fact that $\gamma$ provides a measure of non-equivariance, which we formalize next.
@@
Once again, the proof of a more refined bound can be found in Section A.1. Also note that the bounded assumption on the group action is mild and that in many cases of interest the group action is in fact isometric, i.e., $\left\|\rho_{i}(g)z\right\|=\left\|z\right\|$ (e.g., rotation, translation). Theorem 4.2 shows that by adjusting the degree to which the equivariance constraint in (PI) is relaxed, we can control the extent to which the final solution remains equivariant. While we could replace the equality in (PI) by $|\gamma_{i}|\leq u_{i}$ for $u_{i}\geq 0$, choosing a suitable set of $\{u_{i}\}$ is not trivial, since they must capture how much partial equivariance we expect or can tolerate at each layer. Setting the $u_{i}$ too strictly degrades downstream performance (objective function $\ell_{0}$), whereas setting them loosely interferes with the ability of the model to exploit whatever partial symmetry is in the data. Striking this balance manually is impractical, especially as the number of layers (and consequently $\gamma_{i}$) increases.
Instead, we use the outcomes of Alg. 1 to determine which constraints in (PI) to relax and by how much. Indeed, notice that we only need to relax those $\gamma_{i}$ that do not vanish, seen as they are the ones too strict for the data. We can then leverage the fact that stricter constraints lead to larger $\lambda_{i}$ [53, 57] and relax each layer proportionally to their corresponding dual variable. In doing so, we reduce the number of $\gamma_{i}\neq 0$ as well as their magnitude, thus improving downstream performance (as measured by $\ell_{0}$) while preserving equivariance as much as possible (Theorem 4.2).
This trade-off rationale can be formalized and achieved automatically without repeatedly solving Alg. 1 by using resilient constrained learning [53, 54, 61]. Concretely, we replace (PI) by
@@
where $u\in\mathbb{R}_{+}^{L}$ collects the slacks $u_{i}$ and $\rho>0$ is a proportionality constant (we use $\rho=1$ throughout). Notice that the slacks are not hyperparameters, but optimization variables that explicitly trade off between downstream performance, that improves by relaxing the constraints (i.e., increasing $u_{i}$), and constraint violation (i.e., the level of equivariance). In fact, it can be shown that a solution of (PII) satisfies $u_{i}^{\star}=\lambda^{\star}/\rho$, where $\lambda_{i}^{\star}$ is the Lagrange multiplier of the corresponding constraint [53, Prop. 3]. Using duality, (PII) lead to Alg. 2, where $[z]_{+}=\max(z,0)$ denotes a projection onto the non-negative orthant. Notice that it follows the same steps 3 and 4 of Alg. 1. However, it now contains an additional gradient descent update for the slacks (step 5) and the dual variable updates (step 6) now has a projection due to the constraints in (PII) being inequalities [57]. During training, Alg. 2 therefore automatically balances the level of equivariance imposed on the model ($u_{i}$) and the relative difficulty of imposing this equivariance ($\lambda_{i}$) as opposed to improving downstream performance ($J_{0}$).

### v2

@@ 1 Introduction
Equivariant neural networks (NNs) [Cohen16GroupEqui, pmlr-v70-ravanbakhsh17a, cohen2017steerable, pmlr-v80-kondor18a] leverage known symmetries to improve sample efficiency and generalization, and are widely used in computer vision [Cohen16GroupEqui, cohen2017steerable, Worrall2016HarmonicND, RisiClebschGordan, spherical2018], graph learning [Kip+2017, Gil+2017, Vel+2018, xu2018how], and 3D structure modeling [Sch+2017, Thomas2018TensorFN, Kli+2021, NEURIPS2022_4a36c3c5, Jum+2021, zaverkin:2024]. Despite their advantages, training these networks can be challenging. Indeed, their inherent symmetries give rise to complex loss landscapes [EquivMTL] even when the model matches the underlying data symmetry [EquivMTL, petrache2023approximationgeneralization]. Additionally, real-world datasets frequently include noise, measurement bias, and other symmetry-breaking effects such as dynamical phase transitions or polar fluids [dym2021on, joshiexpressive, Baek_2017, Weidinger2017, Gibb_2024] that undermine the assumptions of strictly equivariant models [findingsymmetrybreaking, approxequiv, symmetrybreaking]. These factors complicate the optimization of equivariant NNs, requiring careful hyperparameter tuning and often resulting in suboptimal performance.
@@
A well-established strategy for approaching difficult optimization problems is to use homotopy (or continuation), i.e., to first solve a simple, surrogate problem and then track its solutions through a sequence of progressively harder problems until reaching the original one [ZissermannHomotopy, YuilleHomotopy]. Continuation has been used for compressive sensing [Osborne2000ANA, Malioutov2005, Donoho2008FastSO] and other non-convex optimization problems [homotopyglobal, iwakiri2022single, xi23homotopy], where it is closely related to simulated annealing strategies [Gelfand90, pmlr-v65-raginsky17a, Belloni2015EscapingTL, OptimSimAnnealing, INGBER199329]. These approaches suggest that introducing equivariance gradually can facilitate training by guiding the model through a series of less stringent intermediate optimization problems. However, implementations of this strategy based on modifying the training objective [EquivMTL] or the underlying model [PennPaper] require domain-specific knowledge to design the intermediate stages and craft loss penalties, limiting practical usability.
@@
This work, therefore, addresses the following question: is there an automated, general-purpose strategy to navigate these intermediate problems that precludes manual tuning and that works effectively for various models and tasks? To this end, we leverage the connection between homotopy approaches and dual methods in optimization, which both provide rigorous frameworks for solving constrained problems by addressing a sequence of relaxed problems with progressively stricter penalties. By casting the training process as a constrained optimization problem, we introduce Adaptive Constrained Equivariance (ACE), a principled framework for automatically relaxing/tightening equivariance in NNs. Our method balances downstream performance (e.g., accuracy) with equivariance, relaxing equivariance at initial stages while gradually transitioning to a strictly equivariant model, as illustrated in Fig.˜1. This transition occurs in an automated, data-driven fashion, eliminating much of the trial-and-error associated with manual tuning.
@@ 2 Related Work
Equivariant NNs encode group symmetries directly in their architecture. In computer vision, for instance, equivariant convolutional NNs accomplish this by tying filter weights for different rotations and reflections [Cohen16GroupEqui, pmlr-v70-ravanbakhsh17a, cohen2017steerable, pmlr-v80-kondor18a]. Equivariant NNs have also been extended to 3D graphs and molecular systems, as is the case of SchNet, that operates directly on atomic coordinates while remaining $\mathsf{E(3)}$‑invariant [Sch+2017], and Tensor Field Networks or $\mathsf{SE(3)}$ Transformers, that extend this to full roto‑translation equivariance using spherical harmonics and attention, respectively [Thomas2018TensorFN, FuchsSE3TF].
Scalable, parameter-efficient architectures that remain equivariant include EGNNs [SatorrasEGNN], SEGNNs [brandstetter2022geometric], and methods such as Vector Neurons [deng2021vn], which lift scalars to 3D vectors, achieving exact $\mathsf{SO}(3)$ equivariance in common layers with minimal extra parameters. Moreover, recent works model 3D trajectories rather than static snapshots, such as EGNO, which uses $\mathsf{SE(3)}$ equivariant layers combined with Fourier temporal convolutions to predict spatial trajectories [xu2024equivariant].
While exact equivariance can improve sample efficiency [GroupThAug, BiettiGroupInvariance, Sannai2019ImprovedGB], it may render the optimization harder, particularly when the data violates symmetry (even if partially) due to noise or other symmetry-breaking phenomena [Baek_2017, Weidinger2017, Gibb_2024]. This has motivated the use of models with approximate or relaxed equivariance based on architectures that soften weight‑sharing, such as Residual Pathway Priors [ResidPathwayPriors] that adds a non‑equivariant branch alongside the equivariant one, or works such as  approxequiv, where weighted filter banks are used to interpolate between equivariant kernels. Recent works have formalized the trade‑offs that appear due to these symmetry violations, establishing generalization bounds that specify when and how much equivariance should be relaxed [petrache2023approximationgeneralization].
Beyond architectural modifications, another line of work relaxes symmetry during training. REMUL [EquivMTL] adds a penalty for equivariance violations to the training objective, adapting the penalty weights throughout training to balance accuracy and equivariance. This procedure does not guarantee that the final solution is equivariant or even the degree of symmetry achieved. An alternative approach modifies the underlying architecture by perturbing each equivariant layer and gradually reducing the perturbation during training according to a user-defined schedule [PennPaper]. This guided relaxation facilitates training and guarantees that the final solution is equivariant, but is sensitive to the schedule. Additional equivariance penalties based on Lie derivative are used to mitigate this issue, increasing the number of hyperparameters and restricting the generality of the method.
@@ 3 Training Equivariant Models
Equivariant NNs improve sample efficiency by encoding transformations from a symmetry group $G$ [GroupThAug, BiettiGroupInvariance, Sannai2019ImprovedGB]. Strictly enforcing equivariance, however, limits the expressiveness of the model and create intricate loss landscapes, which can slow down optimization and hinder their training [petrache2023approximationgeneralization, EquivMTL], particularly in tasks involving partial or imperfect symmetries [dym2021on, joshiexpressive, Baek_2017, Weidinger2017, Gibb_2024].
To temper these trade-offs, recent work relaxes symmetry during training by modifying either the optimization objective or the model itself [PennPaper, EquivMTL, ResidPathwayPriors, approxequiv, petrache2023approximationgeneralization]. Explicitly, let $f_{\theta,\gamma}:X\to Y$ be a NN now parameterized by both $\theta$, the network weights, and $\gamma\in\mathbb{R}^{L}$, a set of homotopic parameters representing the model non-equivariance. The model is therefore equivariant for $\gamma=0$, i.e., $f_{\theta,0}$ satisfies (1), but can behave arbitrarily if $|\gamma_{i}|>0$ for any $i$. Consider training this NN by solving
@@
where $(x_{n},y_{n})$ are samples from a data distribution $\mathfrak{D}$, $\ell_{0}$ denotes the top-line objective function (typically, mean-squared error or cross-entropy), $\ell_{\text{eq}}$ denotes an equivariance metric, and $\alpha,\beta>0$ are hyperparameters controlling their contributions to the training loss. Hence, traditional equivariant NN training amounts to using $\alpha=1$ and $\beta=\gamma=0$ in (2). In [EquivMTL], a non-equivariant model ($\gamma=1$) is trained using $\ell_{\text{eq}}(f)=\sum_{n}\mathbb{E}_{g_{n}}\bigl[\|f\big(\rho_{X}(g_{n})x_{n}\big)-\rho_{Y}(g_{n})f(x_{n})\|^{2}\bigr]$ with $g_{n}$ sampled uniformly from $G$. The values of $\alpha,\beta$ are adapted to balance the contributions of each loss, i.e., to balance accuracy and equivariance. In contrast, [PennPaper] uses an equivariance metric $\ell_{\text{eq}}$ based on Lie derivatives [gruver2023the, otto2024unifiedframeworkenforcediscover], fix $\alpha,\beta$, and manually decrease $\gamma$ throughout training until it vanishes, i.e., until the model is equivariant.
While effective in particular instances, these approaches are based on sensitive hyperparameters, some of which are application-dependent. This is particularly critical in [EquivMTL], where both performance and equivariance depend on the values of $\alpha,\beta$. While [PennPaper] guarantees equivariance by driving $\gamma$ to zero, its performance is substantially impacted by the rate at which it vanishes. It therefore continues to use $\ell_{\text{eq}}$ even though it is not needed to ensure equivariance. What is more, it is once again susceptible to symmetry-breaking effects in the data since it strictly imposes equivariance. In the sequel, we address these challenges by doing away with hand-designed penalties ($\beta=0$) and manually-tuned schedules for $\gamma$. Before proceeding, however, we briefly comment on the implementation of $f_{\theta,\gamma}$.
@@
Throughout this paper, we consider an architecture similar to [PennPaper] built by composing equivariant layers linearly perturbed by non-equivariant components, e.g., a linear layer or a small multilayer perceptron (MLP). Formally, let
@@ 4.1 Strictly equivariant data: Equality constraints
where $(x_{n},y_{n})$ are i.i.d. samples from $\mathfrak{D}$, $\lambda\in\mathbb{R}^{L}$ collects the dual variables [ResilienceConstrained, Chamon20r] $\lambda_{i}$, and $\hat{L}$ is called the empirical Lagrangian [Chamon20p, Chamon23c]. In general, solutions of (PI) and (DI) are not related due to (i) non-convexity, which hinders strong duality [Bertsekas_2009], and (ii) the empirical approximation of the expectation in (PI) by samples. Yet, for rich enough parametrizations $f_{\theta,\gamma}$, constrained learning theory shows that solutions of (DI) do generalize to those of (PI[ResilienceConstrained, Chamon20r, Chamon20p]. In other words, (DI) approximates the solution of the equivariant NN training problem (PI) without constraining the model to be equivariant.
More precisely, (DI) gives rise to a gradient descent (on $\theta,\gamma$) and ascent (on $\lambda$) algorithm (Alg. 1) that acts as an annealing mechanism adapted to the downstream task, taking into account the effect of enforcing equivariance on the objective function. Indeed, by taking $\gamma_{i}^{(0)}=1$, Alg. 1 starts by training a flexible, non-equivariant model (step 3). The gradient step on the homotopy parameters $\gamma_{i}$ (step 4) then simultaneously seeks to reduce the objective function $\ell_{0}$ and bring $\gamma_{i}$ closer to zero, targeting the constraint in (PI). The strength of that bias toward zero (namely, $\lambda_{i}$) depends on the accumulated constraint violations (step 5): the longer $\gamma_{i}$ is strictly positive (negative), the larger the magnitude of $\lambda_{i}$. Hence, the $\gamma_{i}$ can increase to expand the flexibility of the model or decrease to exploit the symmetries in the data. The initial non-equivariant NN is therefore turned into an equivariant one by explicitly taking into account the effects this has on the downstream performance, instead of relying on hand-tuned, ad hoc penalties or schedules as in [EquivMTL, PennPaper].
Naturally, the iterates $\gamma_{i}^{(t)}$ do not vanish in any finite number of iterations. In fact, the behavior of saddle-point algorithms such as Alg. 1 are intricate even in convex settings [approxprimalsolutions, asymp-convergence-primal-dual] and its iterates may, even for small learning rates $\eta_{p},\eta_{d}$, (i) asymptotically approach zero or (ii) display oscillations that do not subside. In case (i), if we stop Alg. 1 at iteration $T$ and deploy its solution with $\gamma=0$ (i.e., $f_{\theta^{(T)},0}$), we introduce an error that depends on the magnitude of $\gamma_{i}^{(T)}$. The following theorem bounds this approximation error for the architecture in (3) under mild assumptions on its components. In particular, note that Ass. 1 holds, e.g., for MLPs with ReLU activations and weight matrices of bounded spectral norms [Horn_Johnson_2018].
@@
A proof of this as well as a more refined bound is available in the Section˜A.6. Theorem 4.1 shows that as long as the $\{\gamma_{i}^{(T)}\}$ are small enough, the error incurred from setting them to zero is negligible. This is not the case, however, when the $\gamma_{i}$ oscillate without subsiding [case (ii)]. As we argue next, this effect can be used to identify and correct for symmetry deviations in the data.
@@
[…]
@@
[…]
@@ 4.2 Partially equivariant data: Resilient constraints
In some practical scenarios, the data used for training may only partially satisfy the equivariance relation (1) due to the presence of noise, measurement bias, or other symmetry-breaking effects [dym2021on, joshiexpressive, Baek_2017, Weidinger2017, Gibb_2024, findingsymmetrybreaking, approxequiv, symmetrybreaking]. In such cases, the equivariance constraints in (PI) become too stringent and the gradient of the objective with respect to $\gamma_{i}$ in step 4 of Alg. 1 does not vanish as $\gamma_{i}$ approaches zero. Alg. 1 will then oscillate as steps 4 and 5 push $\gamma_{i}$ alternatingly closer and further from zero [case (ii)] or $\gamma_{i}$ will settle, but $\lambda_{i}$ will not vanish. These observations can then be used as evidences that the data does not fully adhere to imposed symmetries. This is not necessarily a sign that equivariance should be altogether abandoned, but that we can improve performance by relaxing it. To this end, we rely on the fact that $\gamma$ provides a measure of non-equivariance, which we formalize next.
@@
Once again, the proof of a more refined bound can be found in Section˜A.6. Also note that the bounded assumption on the group action is mild and that in many cases of interest the group action is in fact isometric, i.e., $\left\|\rho_{i}(g)z\right\|=\left\|z\right\|$ (e.g., rotation, translation). Theorem 4.2 shows that by adjusting the degree to which the equivariance constraint in (PI) is relaxed, we can control the extent to which the final solution remains equivariant. While we could replace the equality in (PI) by $|\gamma_{i}|\leq u_{i}$ for $u_{i}\geq 0$, choosing a suitable set of $\{u_{i}\}$ is not trivial, since they must capture how much partial equivariance we expect or can tolerate at each layer. Setting the $u_{i}$ too strictly degrades downstream performance (objective function $\ell_{0}$), whereas setting them loosely interferes with the ability of the model to exploit whatever partial symmetry is in the data. Striking this balance manually is impractical, especially as the number of layers (and consequently $\gamma_{i}$) increases.
Instead, we use the outcomes of Alg. 1 to determine which constraints in (PI) to relax and by how much. Indeed, notice that we only need to relax those $\gamma_{i}$ that do not vanish, seen as they are the ones too strict for the data. We can then leverage the fact that stricter constraints lead to larger $\lambda_{i}$ [ResilienceConstrained, Bertsekas_2009] and relax each layer proportionally to their corresponding dual variable. In doing so, we reduce the number of $\gamma_{i}\neq 0$ as well as their magnitude, thus improving downstream performance (as measured by $\ell_{0}$) while preserving equivariance as much as possible (Theorem 4.2).
This trade-off rationale can be formalized and achieved automatically without repeatedly solving Alg. 1 by using resilient constrained learning [ResilienceConstrained, Chamon20r, Chamon20c]. Concretely, we replace (PI) by
@@
where $u\in\mathbb{R}_{+}^{L}$ collects the slacks $u_{i}$ and $\rho>0$ is a proportionality constant (we use $\rho=1$ throughout). Notice that the slacks are not hyperparameters, but optimization variables that explicitly trade off between downstream performance, that improves by relaxing the constraints (i.e., increasing $u_{i}$), and constraint violation (i.e., the level of equivariance). In fact, it can be shown that a solution of (PII) satisfies $u_{i}^{\star}=\lambda^{\star}/\rho$, where $\lambda_{i}^{\star}$ is the Lagrange multiplier of the corresponding constraint [ResilienceConstrained, Prop. 3]. Using duality, (PII) lead to Alg. 2, where $z]_{+}=\max(z,0)$ denotes a projection onto the non-negative orthant. Notice that it follows the same steps 3 and 4 of Alg. 1. However, it now contains an additional gradient descent update for the slacks (step 5) and the dual variable updates (step 6) now has a projection due to the constraints in (PII) being inequalities [Bertsekas_2009]. During training, Alg. 2 therefore automatically balances the level of equivariance imposed on the model ($u_{i}$) and the relative difficulty of imposing this equivariance ($\lambda_{i}$) as opposed to improving downstream performance ($J_{0}$).